OpenStudy (anonymous):

Find the zeros of the function. f(x) = 9x^3 + 18 x^2 - 135x

OpenStudy (anonymous):

factor out \(x\) then u have \( 9x^3 + 18 x^2 - 135x=x(9x^2+18x-135)=0\) according to the zero product property \(x=0\) or \( 9x^2+18x-135=0\) second is a quadratic equation solve by factoring or quadratic equation formula

OpenStudy (anonymous):

5 and -3

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

simplify quadratic by dividing by 9: x^2 + 2x - 15= 0

OpenStudy (cwrw238):

(x - 3)(x + 5) = 0 x = 3 , -5

OpenStudy (anonymous):

sorry .......... 3 , -5